### v1

@@ B.1 Intermediate sets
Then, we are to generate two distractor sets $E_{\mathrm{n}}$ and $I_{\mathrm{n}}$, representing non-extent object sets and non-intent attribute sets, correspondingly. For each extent $e_{\mathrm{p}}\in E_{\mathrm{p}}$, a distractor object set $e_{\mathrm{n}}$ is generated by randomly selecting $k$ percentage of objects from $e_{p}$ and replacing them with randomly selected objects from the entire object set. In our experiments we set $k$ to $27\%$. If we find this generated $e_{n}$ satisfies that $e_{\mathrm{n}}\notin E_{\mathrm{p}}$ and $e_{\mathrm{n}}\notin E_{\mathrm{n}}$, then we add it into $E_{\mathrm{n}}$ or otherwise we regenerate another $e_{\mathrm{n}}$. Similarly, for each intent $i_{\mathrm{p}}\in I_{\mathrm{p}}$, a distractor attribute set $i_{\mathrm{n}}$ is generated by randomly selecting $k$ attributes from $i_{\mathrm{p}}$ and replacing them with randomly selected attributes from the entire attribute set. If $i_{\mathrm{p}}\notin I_{\mathrm{p}}$ and $i_{\mathrm{p}}\notin I_{\mathrm{n}}$, it is added to $I_{\mathrm{n}}$. The pseudocode of the algorithm is presented in 2.

### v2

@@ B.1 Intermediate sets
Then, we are to generate two distractor sets $E_{\mathrm{n}}$ and $I_{\mathrm{n}}$, representing non-extent object sets and non-intent attribute sets, correspondingly. For each extent $e_{\mathrm{p}}\in E_{\mathrm{p}}$, a distractor object set $e_{\mathrm{n}}$ is generated by randomly selecting $k$ percentage of objects from $e_{p}$ and replacing them with randomly selected objects from the entire object set. If we find this generated $e_{n}$ satisfies that $e_{\mathrm{n}}\notin E_{\mathrm{p}}$ and $e_{\mathrm{n}}\notin E_{\mathrm{n}}$, then we add it into $E_{\mathrm{n}}$ or otherwise we regenerate another $e_{\mathrm{n}}$. Similarly, for each intent $i_{\mathrm{p}}\in I_{\mathrm{p}}$, a distractor attribute set $i_{\mathrm{n}}$ is generated by randomly selecting $k$ attributes from $i_{\mathrm{p}}$ and replacing them with randomly selected attributes from the entire attribute set. If $i_{\mathrm{p}}\notin I_{\mathrm{p}}$ and $i_{\mathrm{p}}\notin I_{\mathrm{n}}$, it is added to $I_{\mathrm{n}}$. The pseudocode of the algorithm is presented in 2.